# Optimisation of the vertex detector and measurement of Higgs decays to second-generation quarks at the CEPC

### v3

@@ 1 Introduction
Electron–positron colliders provide uniquely clean environments that facilitate precision studies of the Higgs boson properties, crucial for probing physics beyond the Standard Model (SM) ES. Among proposed future facilities, the Circular Electron–Positron Collider (CEPC) is designed with a large circumference of 100 km and two interaction points CEPC_CDR_Acc, and will operate at various centre-of-mass energies. According to its Technical Design Report (TDR) CEPC_TDR_Acc; snowmass, the CEPC is planned to accumulate an integrated luminosity of 20 ab-1 at 240 GeV as a Higgs factory an2019precision, yielding $4\times 10^{6}$ Higgs bosons and enabling unparalleled precision in Higgs coupling measurements. In addition, it is capable of running at the Z pole (91.2 GeV), the $W^{+}W^{-}$ threshold (160 GeV), and could potentially be upgraded to reach the $t\bar{t}$ threshold at 360 GeV.
Measurements of the $H\to c\bar{c}$ and $H\to s\bar{s}$ directly test the Yukawa couplings between the Higgs and the second-generation quarks. The SM predicts the $H\to c\bar{c}/s\bar{s}$ has a branching ratio of $2.9\%$$2.3\times 10^{-4}$ SM_BR; Hss_BR; PDG. Notably, these correspond to about 120k $H\to c\bar{c}$ events and about 1k $H\to s\bar{s}$ events at the CEPC when considering all Higgs production modes. Among these, approximately one quarter are produced in the $\nu\bar{\nu}H$ final state from $ZH\,(Z\to\nu\bar{\nu})$ and $WW$-fusion processes. This channel offers a particularly clean environment in which all visible particles originate from the Higgs decay, and therefore provides a convenient benchmark for the simulation study used in this work. With the combination of high-performance detectors and advanced particle reconstruction algorithms, the $H\to c\bar{c}$ is expected to be measured to a percent-level of accuracy, while discovering the $H\to s\bar{s}$ decay mode is a critical and promising possibility at the future Higgs factory.
The vertex detector is essential for Jet Origin Identification (JOI) JOI, which determines the colored parton from which the jet originates, and thus plays a pivotal role in rare Higgs decay measurements. The characterization of the vertex detector largely depends on its geometric configuration, particularly the innermost radius and spatial resolution. In this work, we quantify the relevant detector metrics such as track impact parameter resolution and JOI performance. Furthermore, we evaluate the precision for $H\to c\bar{c}$ and significance for $H\to s\bar{s}$ as functions of these detector parameters, explicitly establishing how detector optimisation translates into physics performance.
@@ 2 CEPC Vertex Detector and Samples
Table 1: The baseline design parameters of the CEPC vertex detector, including position and spatial resolution. CEPC_CDR_Phy
The CEPC baseline detector adopts a Particle Flow Oriented design Ruan_ArborPerf_2018, emphasizing the separation and reconstruction of every final state particle, and measures them in the most suited sub-detectors. A crucial component of the CEPC detector is its vertex system. As referenced in the CEPC Conceptual Design Report (CDR) CEPC_CDR_Phy, the vertex system consists of six concentric cylindrical layers of silicon pixel sensors (Figure 1). The first barrel layer is positioned merely 16 mm from the beam line, and with a per-layer material budget of only $0.15\%\,X_{0}$, it attains a spatial resolution of $2.8\,\mu\text{m}$ for high precision flavour tagging. The beam pipe dominates the upstream material of the vertex detector. In the CEPC TDR design CEPC2025TDR, the Machine Detector Interface(MDI) region is explicitly modelled, including a 0.35 mm beryllium beam pipe with a 10 $\mu m$ inner gold coating and a 0.2 mm cooling-water layer, corresponding in total to about 0.45% $X_{0}$. Each vertex layer contributes approximately 0.06 – 0.28% $X_{0}$, dominated by the sensors and local supports.
@@
The simulated samples correspond to the $\nu\bar{\nu}H$ final state, including both the $ZH$ process with $Z\to\nu\bar{\nu}$ and the $WW$-fusion contribution ($e^{+}e^{-}\to\nu_{e}\bar{\nu}_{e}H$). We use WHIZARD1.95  WHIZARD and Pythia6.4 pythia6 as the event generator and Delphes delphes for the detector simulation. The reconstructed final-state particles are clustered into two jets via the $e^{+}e^{-}$-$k_{t}$ algorithm Suehara:2015ura; Catani:1991hj, and JOI is performed to calculate likelihoods of 11 jet categories for jet-flavour tagging. The Delphes-based results are cross-checked against a Geant4 based full simulation, demonstrating good agreement and robustness. The details can be found in the appendix A.
@@ 3.2 Performance of jet origin identification
The JOI based on the ParticleNet architecture Qu:2019gqs is designed to determine the coloured particle from which a jet originates. For each jet, the algorithm takes the kinematic variables of the reconstructed final-state particles, their particle-type information, the impact-parameter measurements of charged tracks, their energy and momentum distributions, as inputs. It then computes likelihood scores for eleven jet-flavour categories ($b$, $\bar{b}$, $c$, $\bar{c}$, $s$, $\bar{s}$, $u$, $\bar{u}$, $d$, $\bar{d}$, and gluon), which are subsequently employed to construct flavour-tagging observables.
@@
Figure 3 shows the performance of the JOI algorithm. The left-hand panel displays the normalized 11-dimensional jet-flavour migration matrix ($M_{11}$), obtained by assigning each jet to the flavour category with the highest score, thus giving an overview of the tagging and misidentification pattern among all flavours. The right-hand panel provides a complementary flavour-specific view by presenting the $b$-jet tagging performance, obtained from scanning working points with different $b$-jet efficiencies. For each chosen $b$-jet efficiency, the corresponding misidentification rates of the other flavour jets are shown. The $b$-tagging efficiency is defined as the fraction of true $b$-jets whose $b$-score exceeds the threshold, while the mis-ID rate is the fraction of non-$b$ jets passing the same threshold. In the efficiency range of $0.3$ to $0.8$, the JOI achieves misidentification rates that are one to three orders of magnitude lower than those of the multivariable analysis approach based on XGBoost refTDR_report.
@@
Figure 3: Left: The migration matrix of 11-dimensional jet identification of JOI under the CEPC CDR configuration. The matrix is normalized to unity for each truth label (row). Right: The comparison of JOI and XGboost of the $b$-jet tagging efficiency. refTDR_report
@@
From the migration matrices at varying $R_{\mathrm{rad}}$ (see figure 4), we extract key flavour-tagging metrics as shown on the left-hand side of figure 5. As $R_{\mathrm{rad}}$ increases from $0.5$ to $2$, the $b\to c$, $g\to c$ and $c\to s$ misidentification rate rises from about $4\%$ to $7\%$, Meanwhile, the $c\to c$ tagging efficiency drops from $81\%$ to $70\%$, and $s\to s$ from $60\%$ to $54\%$. While the $g\to s$ rate remains stable around $14\%$. The trends indicate that reducing the vertex detector’s innermost radius significantly lowers the $b\to s$ and $c\to s$ misidentification rates, reflecting the crucial role of the vertex in distinguishing heavy-flavour jets from strange jets. In contrast, the impact of vertex detector optimisation on gluon-strange discrimination is limited as gluon jets predominantly originate from the primary vertex, and vertex detector upgrades have little effect on their identification.
@@
Figure 4: The migration matrix of $M_{11}$ with $R_{rad.}$ =0.5 (left) and $R_{rad.}$ =2 (right). The matrix is normalized to unity for each truth label.
@@
with $S_{f\bar{f}}\in[0,1]$ and $\sum_{f}S_{f\bar{f}}=1$. For a given physics measurement, the candidate $H\to f\bar{f}$ signal events are defined as those satisfying $S_{f\bar{f}}$ over the threshold, where the threshold score is scanned to optimise the product $\epsilon\times P$, to balance the signal retention and background suppression, and thus the optimal measurement sensitivity. The efficiency $\epsilon$ is defined as the fraction of true $H\to f\bar{f}$ signal events whose $S_{f\bar{f}}$ value lies above the chosen threshold relative to the total number of simulated signal events. The purity $P$ is the fraction of selected events that truly originate from the $H\to f\bar{f}$ decay, i.e. the ratio of correctly identified signal events to all selected events including background contributions.
Figure 6: Top row: $P$ and $\varepsilon\times P$ as a function of signal efficiency for $H\to c\bar{c}$ at $R_{\rm rad}=0.5$ (left), baseline (middle), and $R_{\rm rad}=2$ (right). Bottom row: the same for $H\to s\bar{s}$.
@@
Using the event-level posterior eq.(4), we derive purity-efficiency curves by varying the score threshold (0 to 1) for $H\to c\bar{c}$ and $H\to s\bar{s}$. As previous results indicate that the impact of spatial resolution is minimal, figure 6 displays only the three representative values of the vertex inner radius. For $H\to c\bar{c}$, the larger branching ratio yields efficiently reconstructed signals, with the maximum of $\varepsilon\times P$ dominated by efficiency. In contrast, $H\to s\bar{s}$ shows steeper purity degradation with tighter cuts, making its maximum of $\varepsilon\times P$ more purity-dependent.
@@
where $s$ and $b$ denote signal and background yields (extracted from score-based selections). The significance follows eq.(6Cowan.
@@
Table 2: Event yields at the optimal cut (maximizing $\varepsilon\times P$) and statistical uncertainty for $H\to c\bar{c}$, normalized to an integrated luminosity of $20\,\text{ab}^{-1}$ in the $\nu\bar{\nu}H$ channel. Columns list signal ($c\bar{c}$), other Higgs decay backgrounds ($b\bar{b}$, $s\bar{s}$, gggg), and relative statistical uncertainty.
@@
Table 3: Event yields at the optimal cut (maximizing $\varepsilon\times P$) and corresponding Asimov significance for $H\to s\bar{s}$, normalized to an integrated luminosity of $20\,\text{ab}^{-1}$in the $\nu\bar{\nu}H$ channel. Columns list signal ($s\bar{s}$), other Higgs decay backgrounds ($b\bar{b}$, $c\bar{c}$, gggg), and significance.
@@ 4 Summary and Discussion
Our observations are consistent with previous studies wu_vertex; Zhu2023ParticleNet, which show that the inner radius of the vertex detector is the most sensitive geometric parameter for flavour tagging and Higgs precision measurements, underscoring its importance in future detector design. The search for $H\to s\bar{s}$ offers a unique probe of second-generation Yukawa couplings and is a major objective for next-generation Higgs factories. Recent global studies project measurement accuracies for this channel in the range of 17% to 140% Zhu:2025eoe; JOI; ECFA, with the spread arising from differences in assumed luminosity, analysis strategies, background modelling, and flavour tagging performance.
@@ Acknowledgements.
We thank Yongfeng Zhu for providing the analysis tools and the useful discussions. This work was supported by National Key Program for S&\&T Research and Development under Contract number 2024YFA1610600, National Key R&\&D Program of China under Contracts No. 2022YFE0116900, the National Key Laboratory of Dark Matter Physics and the National Natural Science Foundation of China under Grant No.W2441004.
@@ Appendix B Simulation Study of the Beam-Induced Background Impact
The influence of beam-induced background on tracking performance has been qualitatively evaluated in the CEPC TDR CEPC2025TDR. As shown in Figure 10, a direct comparison of the impact-parameter resolutions ($d_{0}$ and $z_{0}$) with and without beam-induced background overlay demonstrates no noticeable degradation across the momentum range. Therefore, the pixel size and spatial resolution variations studied in this work remain valid, and the overall conclusions on vertex detector optimization are not affected by beam background effects.

### v4

@@ 1 Introduction
Electron–positron colliders provide uniquely clean environments that facilitate precision studies of the Higgs boson properties, crucial for probing physics beyond the Standard Model (SM) The European Strategy Group (2020). Among proposed future facilities, the Circular Electron–Positron Collider (CEPC) is designed with a large circumference of 100 km and two interaction points Group and Group (2018), and will operate at various centre-of-mass energies. According to its Technical Design Report (TDR) Abdallah and others (2024); Cheng and others (2022), the CEPC is planned to accumulate an integrated luminosity of 20 ab-1 at 240 GeV as a Higgs factory An and others (2019), yielding $4\times 10^{6}$ Higgs bosons and enabling unparalleled precision in Higgs coupling measurements. In addition, it is capable of running at the Z pole (91.2 GeV), the $W^{+}W^{-}$ threshold (160 GeV), and could potentially be upgraded to reach the $t\bar{t}$ threshold at 360 GeV.
[…]
The vertex detector is essential for Jet Origin Identification (JOI) Liang et al. (2024), which determines the colored parton from which the jet originates, and thus plays a pivotal role in rare Higgs decay measurements. The characterization of the vertex detector largely depends on its geometric configuration, particularly the innermost radius and spatial resolution. In this work, we quantify the relevant detector metrics such as track impact parameter resolution and JOI performance. Furthermore, we evaluate the precision for $H\to c\bar{c}$ and significance for $H\to s\bar{s}$ as functions of these detector parameters, explicitly establishing how detector optimisation translates into physics performance.
@@ 2 CEPC Vertex Detector and Samples
Table 1: Baseline design parameters of the CEPC vertex detector. $R$ denotes the radial position of each layer with respect to the beam line, $|z|$ the half length of the layer, $|\cos\theta|$ the corresponding polar-angle acceptance, and $\sigma$ the spatial resolution of the silicon pixel sensors.
The CEPC baseline detector adopts a Particle Flow Oriented design Ruan and others (2018), emphasizing the separation and reconstruction of every final state particle, and measures them in the most suited sub-detectors. A crucial component of the CEPC detector is its vertex system. As referenced in the CEPC Conceptual Design Report (CDR) Group (2018), the vertex system consists of six concentric cylindrical layers of silicon pixel sensors (Figure 1). The first barrel layer is positioned merely 16 mm from the beam line, and with a per-layer material budget of only $0.15\%\,X_{0}$, it attains a spatial resolution of $2.8\,\mu\text{m}$ for high precision flavour tagging. The baseline geometrical parameters and spatial resolutions of the vertex detector layers used in this study are summarised in Table 1. The beam pipe dominates the upstream material of the vertex detector. In the CEPC TDR design Group (2025), the Machine Detector Interface(MDI) region is explicitly modelled, including a 0.35 mm beryllium beam pipe with a 10 $\mu m$ inner gold coating and a 0.2 mm cooling-water layer, corresponding in total to about 0.45% $X_{0}$. Each vertex layer contributes approximately 0.06 – 0.28% $X_{0}$, dominated by the sensors and local supports.
@@
The simulated samples correspond to the $\nu\bar{\nu}H$ final state, including both the $ZH$ process with $Z\to\nu\bar{\nu}$ and the $WW$-fusion contribution ($e^{+}e^{-}\to\nu_{e}\bar{\nu}_{e}H$), corresponding to about $0.9\times 10^{6}$ Higgs bosons at an integrated luminosity of $20~\mathrm{ab}^{-1}$. We use Whizard 1.95 Kilian et al. (2011) and Pythia 6.4 Sjöstrand et al. (2006) as the event generators and Delphes Chen et al. (2017) for the detector simulation. The reconstructed final-state particles are clustered into two jets via the $e^{+}e^{-}$-$k_{t}$ algorithm Suehara and Tanabe (2016); Catani et al. (1991), and JOI is performed to calculate likelihoods of 11 jet categories for jet-flavour tagging. The Delphes-based results are cross-checked against a Geant4 based full simulation, demonstrating good agreement and robustness. The details can be found in the appendix A. An idealised particle identification (PID) model is adopted to study the effects of vertex detector parameters in isolation, without introducing additional dependencies from PID modelling. Studies of the CEPC baseline detector show that kaon identification efficiencies of about 95%–98% can be achieved over a broad kinematic range, approaching the idealised PID assumption Zhu and others (2023); Yu et al. (2025). This supports the use of idealised PID within the conceptual scope of the present work.
@@ 3.2 Performance of jet origin identification
The JOI based on the ParticleNet architecture Qu and Gouskos (2020) is designed to determine the coloured particle from which a jet originates. For each jet, the algorithm takes the kinematic variables of the reconstructed final-state particles, their particle-type information, the impact-parameter measurements of charged tracks, their energy and momentum distributions, as inputs. It then computes likelihood scores for eleven jet-flavour categories ($b$, $\bar{b}$, $c$, $\bar{c}$, $s$, $\bar{s}$, $u$, $\bar{u}$, $d$, $\bar{d}$, and gluon), which are subsequently employed to construct flavour-tagging observables.
@@
Figure 3 shows the performance of the JOI algorithm. The left-hand panel displays the normalized 11-dimensional jet-flavour migration matrix ($M_{11}$), obtained by assigning each jet to the flavour category with the highest score, thus giving an overview of the tagging and misidentification pattern among all flavours. The right-hand panel provides a complementary flavour-specific view by presenting the $c$-jet tagging performance, obtained from scanning working points with different $c$-jet efficiencies. For each chosen $c$-jet efficiency, the corresponding misidentification rates of the other flavour jets are shown. The $c$-tagging efficiency is defined as the fraction of true $c$-jets whose $c$-score exceeds the threshold, while the mis-ID rate is the fraction of non-$c$ jets passing the same threshold. In the efficiency range of $0.2$ to $0.8$, the JOI achieves misidentification rates that are one to three orders of magnitude lower than those of the traditional multivariable analysis approach based on XGBoost Zhang (2025).
@@
Figure 3: Left: The migration matrix of 11-dimensional jet identification of JOI under the CEPC CDR configuration. The matrix is normalized to unity for each truth label (row). Right: The comparison of JOI and XGboost of the $c$-jet tagging efficiency. Zhang (2025)
@@
From the migration matrices at varying $R_{\mathrm{rad}}$ as shown in figure 4, we extract key flavour-tagging metrics as shown on the left-hand side of figure 5. As $R_{\mathrm{rad}}$ increases from $0.5$ to $2$, the $b\to c$, $g\to c$ and $c\to s$ misidentification rate rises from about $4\%$ to $7\%$, Meanwhile, the $c\to c$ tagging efficiency drops from $81\%$ to $70\%$, and $s\to s$ from $60\%$ to $54\%$. While the $g\to s$ rate remains stable around $14\%$. The trends indicate that reducing the vertex detector’s innermost radius significantly lowers the $b\to s$ and $c\to s$ misidentification rates, reflecting the crucial role of the vertex in distinguishing heavy-flavour jets from strange jets. In contrast, the impact of vertex detector optimisation on gluon-strange discrimination is limited as gluon jets predominantly originate from the primary vertex, and vertex detector upgrades have little effect on their identification.
(a) $R_{\mathrm{rad}}=0.5$
(b) $R_{\mathrm{rad}}=2$
@@
Figure 4: The migration matrix of $M_{11}$ with $R_{rad.}$ = 0.5 (a) and $R_{rad.}$ = 2 (b). The matrix is normalized to unity for each truth label.
@@
with $S_{f\bar{f}}\in[0,1]$ and $\sum_{f}S_{f\bar{f}}=1$. The corresponding $S_{ff}$ distributions at event-level are shown in Appendix C. For a given physics measurement, the candidate $H\to f\bar{f}$ signal events are defined as those satisfying $S_{f\bar{f}}$ over the threshold, where the threshold score is scanned to optimise the product $\epsilon\times P$, to balance the signal retention and background suppression, and thus the optimal measurement sensitivity. The efficiency $\epsilon$ is defined as the fraction of true $H\to f\bar{f}$ signal events whose $S_{f\bar{f}}$ value lies above the chosen threshold relative to the total number of simulated signal events. The purity $P$ is the fraction of selected events that truly originate from the $H\to f\bar{f}$ decay, i.e. the ratio of correctly identified signal events to all selected events including background contributions.
Figure 6: Purity $P$, efficiency $\epsilon$, and $\epsilon\times P$ as functions of the discriminant score $S_{ff}$ for $H\to c\bar{c}$ (top) and $H\to s\bar{s}$ (bottom), shown for $R_{\rm rad}=0.5$ (left), baseline (middle), and $R_{\rm rad}=2$ (right). Note that for the $H\to s\bar{s}$, the vertical axis for $\epsilon\times P$ is shown in the range [0, 0.5][0,\,0.5]. The curves are obtained by scanning the $S_{ff}$ selection threshold.
@@
Using the event-level posterior defined in Eq. (4), the purity $P$, efficiency $\varepsilon$, and their product $\varepsilon\times P$ are evaluated as functions of the discriminant score $S_{ff}$. Since previous studies indicate that the impact of the spatial resolution on the flavour-tagging performance is small, figure 6 shows results only for three representative values of the vertex inner radius. For $H\to c\bar{c}$, the large branching ratio yields a high signal rate and good background separation over a broad range of $S_{cc}$. Both the efficiency and purity remain high for moderately tight selections, resulting in a broad maximum of $\varepsilon\times P$. In contrast, for $H\to s\bar{s}$, the maximum of $\varepsilon\times P$ is reached at relatively loose selections. Although the purity improves with increasing $S_{ss}$, the rapid loss of signal efficiency dominates the behaviour of $\varepsilon\times P$, making the optimal working point primarily driven by signal efficiency.
@@
where $s$ and $b$ denote signal and background yields (extracted from score-based selections). The significance follows eq.(6Cowan et al. (2011).
@@
Table 2: Event yields at the optimal cut (maximizing $\varepsilon\times P$) and statistical uncertainty for $H\to c\bar{c}$, normalized to an integrated luminosity of $20\,\mathrm{ab}^{-1}$ in the $\nu\bar{\nu}H$ channel, corresponding to about $0.9\times 10^{6}$ Higgs bosons. Columns list the signal ($c\bar{c}$), other Higgs decay backgrounds ($b\bar{b}$, $s\bar{s}$, gggg), and the relative statistical uncertainty.
@@
Table 3: Event yields at the optimal cut (maximizing $\varepsilon\times P$) and corresponding Asimov significance for $H\to s\bar{s}$, normalized to an integrated luminosity of $20\,\text{ab}^{-1}$ in the $\nu\bar{\nu}H$ channel, corresponding to about $0.9\times 10^{6}$ Higgs bosons. Columns list signal ($s\bar{s}$), other Higgs decay backgrounds ($b\bar{b}$, $c\bar{c}$, gggg), and significance.
@@ 4 Summary and Discussion
Our observations are consistent with previous studies Wu et al. (2018); Zhu et al. (2023), which show that the inner radius of the vertex detector is the most sensitive geometric parameter for flavour tagging and Higgs precision measurements, underscoring its importance in future detector design. The search for $H\to s\bar{s}$ offers a unique probe of second-generation Yukawa couplings and is a major objective for next-generation Higgs factories. Recent global studies project measurement accuracies for this channel in the range of 17% to 140% Zhu et al. (2025); Liang et al. (2024); Altmann and others (2025), with the spread arising from differences in assumed luminosity, analysis strategies, background modelling, and flavour tagging performance.
@@ Acknowledgements.
We thank Yongfeng Zhu for providing the analysis tools and the useful discussions. This work was supported by National Key Program for S&\&T Research and Development under Contract No. 2024YFA1610603, 2024YFA1610600, National Key R&\&D Program of China under Contracts No. 2022YFE0116900, the National Key Laboratory of Dark Matter Physics and the National Natural Science Foundation of China under Grant No.W2441004.
@@ Appendix B Simulation Study of the Beam-Induced Background Impact
The influence of beam-induced background on tracking performance has been qualitatively evaluated in the CEPC TDR Group (2025). As shown in Figure 10, a direct comparison of the impact-parameter resolutions ($d_{0}$ and $z_{0}$) with and without beam-induced background overlay demonstrates no noticeable degradation across the momentum range. Therefore, the pixel size and spatial resolution variations studied in this work remain valid, and the overall conclusions on vertex detector optimization are not affected by beam background effects.
@@
Appendix C Event-level distributions of the discriminant $S_{ff}$
To provide a more intuitive interpretation of the efficiency–purity curves shown in Figure 6, we present in Figure 11 the distributions of the event-level discriminant $S_{ff}$ for the signal and the main Higgs decay backgrounds. The horizontal axis corresponds to the value of $S_{ff}$, while the vertical axis shows the number of events per bin. As discussed in the main text, varying the selection on $S_{ff}$ corresponds to integrating these distributions above the chosen threshold. Tightening the selection requirement progressively shifts the selected events towards the right-hand side, while looser selections include an increasing fraction of events at small $S_{ff}$ values.
Figure 11: Distributions of the event-level discriminant $S_{ff}$ for the baseline vertex detector configuration. Left: $S_{cc}$ distribution for the $H\to c\bar{c}$. Right: $S_{ss}$ distribution for the $H\to s\bar{s}$. The signal and the main Higgs decay backgrounds are shown separately. The vertical axis indicates the number of events per bin on a logarithmic scale.